Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, ruggi, vardaro - 0:34 Friday 14 January 2022 (54476) Print this report
9 hours and 45 min of lock: SRCL longitudinal offset scan, SRCL, and DARM boost, 1f automation test

The goals of the shift were:

  1. Find an angular position of SR with high optical spring, low dark fringe power and high 56MHz sidebands power and change the SRCL loop longitudinal set point in order to remove the optical spring
  2. Add a boost in the DARM and SRCL filters
  3. Test the automation of the SSFS engagement + switch of SRCL, PRCL and MICH to 1f signals
  4. Put in the automation the engagement of the arm cavities dithering

We started the shift with the Interferometer locked since 13:01 UTC and we unlocked it by hand at 21:46:00 UTC in order to test the automation improvement. Figure 1 shows the trend of the long lock. Observing B7 and B8 DC we can see that they are diverging.

Part 1: Longitudinal offset of the signal recycling

At the beginning of the shift the SR angular working point was set in order to minimize the optical spring. At 16:35 UTC we moved SR_MAR_TY by 1urad in order to darken the fringe. Then we scanned the longitudinal set point with the aim to minimize again the optical spring:

  • 16:59:00 SRCL SET = 0
  • 17:01:20 SRCL SET = 0.5
  • 17:07:020 SRCL SET = 2.0
  • 17:26:00 SRCL SET = 3.0
  • 17:34:30 SRCL SET = 4.0
  • 17:40:15 SRCL SET = 5.0
  • 17:44:50 SRCL SET = 6.0
  • 17:52:40 SRCL SET = 6.5

Increasing the SRCL longitudinal set point we could lower the optical spring pole up to quite zero maintaining a good dark fringe level and a good stability of the ITF. We left this longitudinal set point up to the end of the lock.

Figure 2 shows the DARM Open Loop transfer function during the SRCL longitudinal scan. Figure 3 shows the transfer function between DARM and B7 DC during the scan, we can see a good coherence between the two signals and that the magnitude decreased during the scan.

Figure 4 shows the average between 30 and 70 Hz of the Transfer function between DARM and B7_DC as function of the SRCL longitudinal set point. We can see a linear behaviour and that this is a good figure of merit for a fast estimation of the optical spring.

After the SRCL SET SCAN we moved the BS MAR TY in order to darken the fringe at 18:04 UTC we scanned also MAR TX but was already in a good position. After this angular adjustment we did not move the ITF working point up to the end of the lock.

 

Part 2: DARM and SRCL boost

  • 18:12:28 UTC we moved the SRCL UGF servo set point from 10 to 15
  • 18:31:40 UTC we switched off the DARM noise (it was on since the beginning of the shift)
  • 18:37:30 UTC we injected noise in SRCL with " LSC_noise_MICHband" filter and A= 5e-2
  • 18:43:40 UTC we injected noise in MICH with " LSC_noise_MICHband" filter and A= 4e-2
  • 19:05:00 we reduced the lines of the longitudinal DOF servo UGF loop and we took clean data

We modified the SRCL loop adding a boost (pole 1.5Hz Q=1, zero 5Hz Q=0.8) and we moved the pole at 50 Hz tp 60 Hz, the zero at 80 Hz to 90Hz and the pole at 70 Hz to 80 Hz. The new filter is saved as "SRCL_ctrl" and it replace "SRCL_ctrl"

We modified the DARM loop adding a boost (pole 0.5 Q=1.2, zero 12 Q=1, Gain 1 @ 20Hz ). The boost is saved in the "Arm_boost2" filter and should be multiplied by the current filter "DARM_CARM0_ctrl"

  • 21:24:10 UTC we engaged the new SRCL filter
  • 21:20:30 UTC we injected noise in SRCL
  • 21:34:10 UTC we engaged the DARM boost
  • 21:41:10 UTC we injected noise in DARM

Figures 5 - 6 - 7 - 8 - 9 - 10 show the simulation of the SRCL loop. Figure 11 shows the LSC_SRCL FFT with (purple) and without (blue) the new filter

Figures 12 - 13- 14 - 15 - 16 - 17 show the simulation of the DARM loop. Figure 18 shows the LSC_DARM FFT with (purple) and without (blue) the new filter.

In figure 19 the DARM error signal before and after the engagement of the new filter.

 

Part 3. Work on the automation

  • We restarted tango SAT Server from Astor to load the new gname added by Paolo {WE, WI, NE, NI} MIR_{X,Y}_SET 
  • We restarted the PySat Process from VPM
  •  We restored the set point to 0 in the ITF_LOCK DOWN state
  • We tested the engagement of the SSFS and the hand off of MICH, PRCL and SRCL to 1f signals in CARM null. The new state is CARM_NULL_1f.
  • We wrote the dither_carm_null python Funcrion to engage the dithering loop of the Arms in CARM null 1f. We saved it in the ITF_lock_library_HP.py.
    • The function close the dithering of all the horizontal loops and the vertical loops only for the north arm
    • we found an offset on WE Y but we expected no offset on them. We left it
    • dither_carm_null(0) to open the dithering
    • dither_carm_null(1) to close the dithering

N. B. the function for the dithering is not tested

Close and reopen the Metatron ITF_Lock GUI

During the test of 1f hand-off we could reach CARM null state twice without problem.

Figure 20 the new layout of ITF lock node.

We left the Arms locked on the IR

Images attached to this report
Non-image files attached to this report
Comments to this report:
flaminio - 8:52 Friday 14 January 2022 (54478) Print this report
Nice job!
Search Help
×

Warning

×